OpenStudy (anonymous):

if a number decreases from 400 to 94 what is the rate of decrease?

OpenStudy (anonymous):

Do you mean the % decrease?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

% decrease = (decrease / original) * 100 decrease = 400 - 94 = 306 original = 400 So, % decrease = (306 / 400) * 100